OpenStudy (anonymous):

7(5x - 2) = 6(6x - 1) x =

OpenStudy (anonymous):

well expand the brackets and show me wat u get :)

OpenStudy (anonymous):

-2 on both sides?

OpenStudy (anonymous):

no, ur supposed to expand the brackets and make x the subject of the formula :)

OpenStudy (anonymous):

7(5x - 2) = 6(6x - 1) 35x - 14 = 36x - 6 continue from here :)

OpenStudy (anonymous):

please kill me

OpenStudy (anonymous):

no! D: humanity needs ur looks to be passed down lol

OpenStudy (anonymous):

-8 = x cool? :)

OpenStudy (anonymous):

awwww haha thank you

OpenStudy (anonymous):

very cool

OpenStudy (anonymous):

:)